What mass of neon gas is required to fill a 4.37-L container to a pressure of 1.28 atm at 27°C?
Answer #1

I= (27 + 273 15 )k Ti 300.15k v=4.37L p=1.98 atm R = 0.0821 L.atm. mot! k-1 molar man of Ne m= 20.17979 Ideal ger equation PV